OpenStudy (anonymous):

Solve x^2 + 7x = 1 this is quadratics btw

OpenStudy (callisto):

x^2 + 7x = 1 x^2 + 7x -1= 0 Use the quadratic formula again \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] Put a =1, b=7 and c=-1 into the equation, what is(are) the answer(s) you get?

OpenStudy (anonymous):

x equals the quantity of negative 7 plus or minus the square root of 45 all over 2 ?

OpenStudy (callisto):

Not really.... Value of square root is not correct :|

OpenStudy (anonymous):

it should be 53 under the square root since it's 49+4

OpenStudy (anonymous):

Callisto may I ask how u entered the fraction for quadratic formula in here, I can't seem to find the fraction symbol in the Equation tab?

OpenStudy (callisto):

\frac{numerator}{denominator}